OpenStudy (anonymous):

Let W be a random variable giving the number of heads minus the number of tails in three tosses of a coin. (i) List the elements of the sample space S for the three tosses of the coin and to each sample point assign a value w for W. (ii) Find the probability distribution of the random variable W assuming that the coin is biased so that a head is twice as likely to occur as a tail

OpenStudy (zarkon):

do you know what \(S\) is?

OpenStudy (paxpolaris):

(i) \[\begin{matrix}\text{Sample} & W \\ HHH & 3 \\ HHT & 1 \\ HTH & 1 \\ HTT & -1 \\ THH & 1 \\ THT & -1 \\ TTH & -1 \\ TTT & -3\end{matrix}\]

OpenStudy (paxpolaris):

i suppose you were looking for something like that...

OpenStudy (paxpolaris):

(ii) hint: a sequence with 1H is twice as likely than the sequence with no H (TTT). a sequence with 2H is 2*2=4times as likely than the sequence with no H. the sequence with 3H is 2*2*2=8times as likely than the sequence with no H.